OpenStudy (anonymous):

Find the volume of the graph of x=sqrt(1+x) from y=-1 to 3 rotated around the y-axis?

OpenStudy (amistre64):

might have to draw a graph to see what going on

OpenStudy (anonymous):

Yeah...I'm not sure how to do that on here...I didn't think the question made much sense when I typed it like that.

OpenStudy (amistre64):

x = sqrt(1+x) ? or is that a y in there someplace?

OpenStudy (anonymous):

yes, sorry..it's x=sqrt(1+y).

OpenStudy (amistre64):

which would be easier for this; shells or discs?

OpenStudy (anonymous):

discs

OpenStudy (amistre64):

this is the same result as if we were doing: y = sqrt(x+1) from x=[-1,3] around the x axis right?

OpenStudy (anonymous):

that is it, but it is being rotated around the y axis

OpenStudy (amistre64):

the equation is the same regardless of where you put the variable in this case; the same shape results; but ill stick with the original all the same

OpenStudy (amistre64):

now, what we want to determine is the area of a typical circle and then add them all up from -1 to 3

OpenStudy (amistre64):

the area of a circle is pi r^2 right?

OpenStudy (anonymous):

yeah...I have the equation as being \[\pi \int\limits_{-1}^{3}(\sqrt{y+1}^{2}dy?\]

OpenStudy (amistre64):

thats correct; now the ^2 and the sqrt "cancel" which leaves us with: pi. {S} (y+1) dy ; [-1,3] right?

OpenStudy (anonymous):

yeah..and I solved that for what I thought was pi(y^2/2 + y) and from [-1,3] I thought it was 12 pi, but that is incorrect.

OpenStudy (amistre64):

lets see .... pi (9/2+3) - pi (1/2 - 1) pi (9/2 +3 -1/2 +1) pi (8/2 + 4) ... you see where this is headed?

OpenStudy (amistre64):

being able to integrate is no good if you forget how to do the simple math afterwards ;)

OpenStudy (amistre64):

we can save ourselves some mathical skills by lifting the whole thing up by 1 pi. {S} y dy ; [0,4], the 0 is useless ... pi y^2/2 = pi 16/2 = 8

OpenStudy (amistre64):

8 pi ....

OpenStudy (anonymous):

I see what I did wrong...I made the first term 9/2 + 9 because I squared the second y...silly mistake

OpenStudy (anonymous):

I like the idea of lifting it all up by one...hadn't though of that one. Thanks for the help.

OpenStudy (amistre64):

youre welcome :)